# APEX-MR: Multi-Robot Asynchronous Planning and Execution for Cooperative Assembly

### v1

@@ Abstract
Compared to a single-robot workstation, a multi-robot system offers several advantages: 1) it expands the system’s workspace, 2) improves task efficiency, and more importantly, 3) enables robots to achieve significantly more complex and dexterous tasks, such as cooperative assembly. However, coordinating the tasks and motions of multiple robots is challenging due to issues, e.g., system uncertainty, task efficiency, algorithm scalability, and safety concerns. To address these challenges, this paper studies multi-robot coordination and proposes APEX-MR, an asynchronous planning and execution framework designed to safely and efficiently coordinate multiple robots to achieve cooperative assembly, e.g., LEGO assembly. In particular, APEX-MR provides a systematic approach to post-process multi-robot tasks and motion plans to enable robust asynchronous execution under uncertainty. Experimental results demonstrate that APEX-MR can significantly speed up the execution time of many long-horizon LEGO assembly tasks by 48% compared to sequential planning and 36% compared to synchronous planning on average. To further demonstrate the performance, we deploy APEX-MR to a dual-arm system to perform physical LEGO assembly. To our knowledge, this is the first robotic system capable of performing customized LEGO assembly using commercial LEGO bricks. The experiment results demonstrate that the dual-arm system, with APEX-MR, can safely coordinate robot motions, efficiently collaborate, and construct complex LEGO structures. Our project website is available at https://intelligent-control-lab.github.io/APEX-MR/.
@@ I Introduction
Multi-robot manipulation is critical in robotic applications, such as industrial assembly [37, 3], material handling [50], and object arrangement [12, 14], etc. Compared to a single-robot setup, a multi-robot arm system can easily expand the system’s overall reachable area. Besides, with a team of robot arms, a task can be accomplished more efficiently [49] by having each robot execute individual tasks parallelly. In addition to improving task efficiency, multi-robot can achieve significantly greater dexterity and is necessary in many applications that require collaborations, e.g., cooperative assembly, since certain tasks cannot be done with only one arm.
@@
LEGO assembly is an example of cooperative assembly. In a LEGO structure, bricks are assembled by forcing the top knobs of a brick into the bottom cavities of another brick. The bricks are held together passively, i.e., by the friction in the knob-to-cavity connections. Thus, the connections are not rigid, and subsequent manipulation, if performed inappropriately, can easily break the existing structure. Due to the nature of the passive connection, a multi-robot system is necessary for such cooperative assembly. Fig. 1 demonstrates example assembly operations in constructing a LEGO structure. Fig. 1(a) shows the robot picking up a brick by disassembling it from the LEGO plate (i.e., pick). Fig. 1(b) illustrates the robot assembling a brick by placing it at the desired location and forcing a solid connection (i.e., place-down). One robot is sufficient for these two tasks since the manipulated object is fully supported and the operation would not collapse the existing structure. On the other hand, Fig. 1(c)-(e) showcase operations that require multi-arm collaboration. In Fig. 1(c), the upper robot is assembling a LEGO brick on top of the character ‘S’. To establish a solid connection, the robot needs to press down and force the knob insertion due to the passive connection nature. However, since the existing connections are non-rigid, the place-down operation would break the existing overhanging structure. Therefore, the lower robot is necessary to support and stabilize the structure from below (i.e., support-bottom). Similarly in Fig. 1(e), the lower robot assembles a brick from the bottom onto the character ‘S’ by pushing it up and forcing the connection (i.e., place-up). The place-up operation would break the existing structure, and thus, the upper robot is required to press down (i.e., support-top) in order to stabilize the structure. In addition to cooperative assembling, the multi-robot collaboration also enables more dexterous object manipulation, e.g., reorienting bricks in hand. As shown in Fig. 1(d), the upper robot grabs a brick from its top initially. To have the robot grab the brick from its bottom, we can have the upper robot handover the brick to the lower robot. With the capability of reorienting in-hand objects, the robot can subsequently accomplish the assembly from the bottom as illustrated in Fig. 1(e). Despite being a toy brand, LEGO has been widely used in entertainment, education, prototyping, etc. It is ideal for assembly benchmarking since it is a low-cost, standardized, and highly customizable assembly platform. Meanwhile, constructing LEGO structures is a challenging contact-rich manipulation problem due to the high-precision requirement and non-rigid connections. Thus, we use LEGO as our cooperative assembly benchmarking platform and the remaining discussion is in the context of LEGO assembly.
@@
Coordinating the tasks and motions of multiple robots to accomplish cooperative assembly is challenging for several reasons. First, a system with more than a single robot introduces overhead and algorithmic challenges in their coordination. As shown in Fig. 1 certain operations involve contacts with objects (i.e., pick, place-down, place-up, handover, support-bottom, and support-top) while some do not (i.e., transit and wait). There exist delays in controlling the robot and receiving sensor feedback, or even contingencies due to unexpected events when performing contact-rich operations, which would cause delays or stops to one or more robots. The uncertainty makes it challenging to safely coordinate the robots throughout the assembly process. Second, due to the collaboration need, robots often need to operate closely to each other as illustrated in Fig. 1(c)-(e). Even when executing under uncertainty in a confined shared workspace, robots must always avoid collisions. Third, it is desired that the task can be accomplished more efficiently by a multi-robot system. Thus, robots must have a comprehensive understanding of the cooperative assembly task and plan optimized motions to reduce the completion time, instead of frequently stopping to avoid collisions. Lastly, the multi-robot system should scale as the complexity of assembly design increases. It is necessary that the algorithm can scale to larger and more complex tasks.
@@
Many recent works have proposed methods for planning multiple robot arms for assembly [22, 5], rearrangement [12, 13], or general manipulation tasks [45]. However, these planners often assume a sequential execution order and synchronously moving robots, where all robots start a task at the same time and wait for other robots to finish. We take a different lens and propose APEX-MR, an asynchronous planning and execution framework, to robustly, efficiently, and safely coordinate multiple robots. Given an assembly task sequence, APEX-MR leverages integer-linear programming (ILP) to distribute the tasks to the robots and generate a sequential robot plan. Most importantly, APEX-MR extends the temporal plan graph (TPG) [20] to multi-robot-arm systems and post-processes the sequential robot plan for asynchronous execution. Specifically, the TPG captures all dependencies between different robots’ tasks and motions and generates a partial-order graph, which can significantly reduce unnecessary wait time and is robust to execution delay and uncertainty. To highlight the applicability of our proposed algorithm within a full multi-robot assembly pipeline, we deploy the proposed APEX-MR to a dual-industrial-arm system to construct complex customized LEGO structures up to 258 objects in simulation and up to 47 objects in the real world. In summary, our contributions are as follows:
@@
The rest of this paper is organized as follows: Sec. II discusses relevant works. Sec. III deliberates the input that APEX-MR consumes and assumptions we have. Sec. IV introduces our proposed APEX-MR, an asynchronous planning and execution framework for multi-robot cooperative assembly. Sec. V shows the experiment results and demonstrates the deployment to a bimanual system for LEGO assembly. Sec. VI discusses limitations and future works. Sec. VII concludes the paper.
@@ II Related Works
Multi-Agent Path Finding and Execution Multi-agent path finding (MAPF) [54] studies how to coordinate a large team of mobile robots in a discretized world environment, often on 2D grid worlds representing warehouse settings [32] or pre-defined roadmaps [21]. State-of-the-art MAPF algorithms can plan near-optimal collision-free paths for hundreds of robots in seconds [26].
@@
However, these MAPF algorithms achieve such impressive efficiency at the cost of neglecting robot kinematics and execution uncertainty. As a result, there is growing research within the MAPF community focused on the efficient and robust execution of (imperfect) MAPF plans on real robots. One of the most widely adopted frameworks is the temporal plan graph (TPG) [20], originally proposed to post-process MAPF plans to meet robot kinematics by enforcing passing orders at locations visited by multiple robots. TPG has since been extended to execution frameworks for MAPF under uncertainty [33] and mobile robot coordination in warehouses [19, 59]. Recent advancements introduce bidirectional TPG [56] and switchable TPG [1, 23], which further enhance TPG by allowing flexible passing orders at certain locations.
@@
Given the success of TPG in coordinating mobile robots, we aim to explore its applicability in coordinating robotic arms. A key distinction of traditional TPG versus our use case is that the robot kinematic is more complex, and may change over time as the robot arm picks up different object, which is a key focus in this paper.
@@
Multi-Robot Arm Motion Planning Motivated by the rise of bimanual manipulation systems, many early works in the field study the problem of dual arm motion planning [51]. One naive approach that scales single-robot motion planning methods to two or more robots is to plan in the composite joint space, with methods such as RRT-Connect [25], BIT* [11], or graph of convex sets [35]. However, due to the curse of dimensionality, these methods struggle to find high-quality solutions as the number of robots increases. Other common strategies include building a composite roadmap from the Cartesian product of individual roadmaps [15], coordinating the speed of individually planned motions [2], or using prioritized planning [8]. More recently, more specialized multi-robot motion planners have been proposed that are based on roadmaps (dRRT [53], dRRT* [48], and CBS-MP [52]) or utilizing multi-agent path finding techniques [44, 45]. Some approaches also use online planning or control techniques to generate motions in real-time. Zhang and Pecora [61] proposes an online motion coordination technique, in which they plan all robot’s paths offline and a pairwise collision matrix between robots. Then, the speed of each robot can be efficiently planned online to avoid collisions, even in the presence of execution delays. However, their method cannot always find a feasible motion and relies on a task reallocation process to avoid deadlock. Other techniques such as a distributed model predictive controller [10] or a dynamical system approach [38] can also be used to generate multi-robot arm motions in real-time, but are not tailored towards long-horizon planning tasks.
@@
Multi-Robot Task and Motion Planning Beyond motion planning, multi-robot arm task and motion planning (MR-TAMP) have been studied since the 1990s, e.g., object pick and place [24], and many of which are designed for a dual arm setting [17, 12, 14, 51]. Similar to our method, these methods take a two-staged approach in which they first generate a task plan with robot assignment and grasp poses, then search for corresponding motion plans with composite-state-space or prioritized planning. The authors of dRRT* have proposed extending dRRT* to multi-modal roadmaps with given possible pick-up and hand-off configurations and searching for a task and motion plan directly [47, 46]. Given that motion planning calls tend to be more costly, another approach is to generate promising task plans first in a lazy manner and subsequently find corresponding motion plans and backtrack when required. This approach can be implemented with a greedy method [18], with a mixed-integer linear program [49, 5, 31] or through a satisfiability modulo theories solver [42]. However, a major drawback is that task planning often assumes motions to be synchronous, which can be suboptimal in practice. Our TPG framework is designed to be complementary to these synchronous tasks and motion planners, as it can relax the synchronicity assumption with post-processing and shorten the makespan of the overall plan. Another significant concern is that many MR-TAMP methods are designed for simple environments such as object pick and place and planned robot trajectories are executed in open-loop or only in simulation. Our framework scales to more complex environments and is designed to integrate with more challenging manipulation skills such as a force-feedbacked controller.
@@
Multi-Robot Arm Motion Execution Most existing work executes their multi-robot on real robots in a synchronized way. Since the popular motion planning framework MoveIt [7] does not natively support moving multiple robot arms asynchronously, some recent work have sought to address this and enable asynchronous execution. Meehan et al. [36] adds a trajectory reservation component and treats the entire trajectory of a moving robot arm as a static obstacle when planning another arm’s motion, which can be too conservative and cause deadlock. Stoop et al. [55] uses a central scheduler to check if a new trajectory collides with previously scheduled trajectories and executes it asynchronously if there is no collision. Otherwise, the new trajectory waits for the conflicting previous trajectory to finish before it can start. However, their methods do not account for execution delays, may wait longer than necessary, and directly modify the MoveIt! software stack. In contrast, our TPG formulation is robust to arbitrary delays by design, minimizes robot wait time, and can be implemented without modification to MoveIt.
@@
Robotic LEGO Assembly Automating LEGO assembly using robots is challenging due to the high-precision requirement, tiny sizes of LEGO bricks, and non-rigid connections in the structure. Most of the existing works address the LEGO assembly problem in simulation [43, 39], which cannot be generalized to physical assembly due to the lack of simulators to simulate the connections between LEGO bricks. Recent works [9, 16, 27, 30] assume the structure is fully supported and only consider placing bricks on top of others, which are limited to assembling simple structures. [34] considers assembly using customized brick toys, which does not apply to LEGO assembly. In this paper, we apply APEX-MR to a bimanual system to construct complex customized LEGO structures beyond simple stacking.
@@ III Preliminaries
To coordinate a multi-robot system to perform cooperative assembly tasks, we assume three inputs are provided.
@@
Environment Setup We assume the environment setup, including (a) geometries of all robots, (b) poses of objects to be manipulated $\mathcal{B}=[b_{1},b_{2},\dots,b_{N_{b}}]$, and (c) states of all obstacles, is known as shown in the input section in Fig. 2. We assume the system consists of $N$ robots, and $N_{b}$ is the number of objects that can be used for the assembly. An object $b_{i}$ is semi-static, meaning it can be grasped, attached, and moved by the robots. Note that since duplicate objects are common in assemblies, $b_{i}$ and $b_{j}$ can be identical objects, e.g., identical $2\times 2$ bricks in the character ‘S’ shown in Fig. 2.
@@
Assembly Plan Given an assembly design with $N_{a}$ objects, we assume the assembly plan $A=[a_{1},a_{2},\dots,a_{N_{a}}]$ is provided as shown in the input section in Fig. 2. Each step $a_{j}$ refers to an object, such as a $1\times 2$ brick. The assembly plan specifies the order in which each object should be assembled. Note that due to duplicate objects in $\mathcal{B}$, there are multiple candidates $b_{i},\dots,b_{k}$ that can be used to accomplish an assembly step $a_{j}$ in $A$. Furthermore, we require the assembly plan $A$ to be physically valid [58, 28], ensuring that each task can be performed in reality. Specifically, the partially assembled structure after each step $a_{j}$ must be physically stable, and there exists at least one feasible grasp pose for each step. The assembly sequence also specifies whether each step $a_{j}$ requires two robots for cooperative assembly and the specific type of cooperation needed (i.e., support or reorientation). Lastly, we assume $\mathcal{B}$ is sufficient so that each object required for $a_{j}$ can be found in the environment.
@@
Manipulation Skills We assume that the robot manipulation skills are predefined and known. We denote the skill set as $\mathcal{S}=[s_{1},s_{2},\dots,s_{N_{s}}]$, where $N_{s}$ is the number of skills an individual robot has. For instance, a skill can be, inserting a pin, fastening a screw, picking up an object, etc. In our case of LEGO assembly, each skill $s_{i}$ is a composite of multiple motions parametrized by the pose of the manipulated object, the robot end-effector, and parameters learned from demonstrations. We assume an algorithm, such as interpolation or RRT-Connect, can generate a reference robot trajectory for the purpose of computing a collision-free coordination schedule. During execution, each skill is executed with a feedback controller so that its exact execution time is unpredictable. Example manipulation skills are shown in Fig. 1, and more details are discussed in Sec. V.
@@
In addition to these inputs, we assume the existence of a collision-free HOME pose for each robot that never blocks other moving robots from executing their tasks. Robots can also transit between different poses, or wait and hold their current pose in place.

### v2

@@ Abstract
Compared to a single-robot workstation, a multi-robot system offers several advantages: 1) it expands the system’s workspace, 2) improves task efficiency, and, more importantly, 3) enables robots to achieve significantly more complex and dexterous tasks, such as cooperative assembly. However, coordinating the tasks and motions of multiple robots is challenging due to issues, e.g., system uncertainty, task efficiency, algorithm scalability, and safety concerns. To address these challenges, this paper studies multi-robot coordination and proposes APEX-MR, an asynchronous planning and execution framework designed to safely and efficiently coordinate multiple robots to achieve cooperative assembly, e.g., LEGO assembly. In particular, APEX-MR provides a systematic approach to post-process multi-robot tasks and motion plans to enable robust asynchronous execution under uncertainty. Experimental results demonstrate that APEX-MR can significantly speed up the execution time of many long-horizon LEGO assembly tasks by 48% compared to sequential planning and 36% compared to synchronous planning on average. To further demonstrate performance, we deploy APEX-MR in a dual-arm system to perform physical LEGO assembly. To our knowledge, this is the first robotic system capable of performing customized LEGO assembly using commercial LEGO bricks. The experimental results demonstrate that the dual-arm system, with APEX-MR, can safely coordinate robot motions, efficiently collaborate, and construct complex LEGO structures. Our project website is available at https://intelligent-control-lab.github.io/APEX-MR/.
@@ I Introduction
Multi-robot manipulation is critical in robotic applications, such as industrial assembly [37, 3], material handling [50], and object arrangement [12, 13], etc. Compared to a single-robot setup, a multi-robot arm system can easily expand the system’s overall reachable area. Besides, with a team of robot arms, a task can be accomplished more efficiently [49] by having each robot execute individual tasks in parallel. In addition to improving task efficiency, multi-robot systems can achieve significantly greater dexterity and are necessary in many applications that require collaborations, e.g., cooperative assembly, since certain tasks cannot be done with only one arm.
@@
LEGO assembly is an example of cooperative assembly. In a LEGO structure, the bricks are assembled by forcing the top knobs of a brick into the bottom cavities of another brick. The bricks are held together passively, i.e., by the friction in the knob-to-cavity connections. Thus, the connections are not rigid and subsequent manipulation, if performed inappropriately, can easily break the existing structure. Due to the nature of the passive connection, a multi-robot system is necessary for such cooperative assembly. Fig. 1 demonstrates example assembly operations in the construction of a LEGO structure. Fig. 1 (a) shows the robot picking up a brick by disassembling it from the LEGO plate (i.e., pick). Fig. 1 (b) illustrates the robot assembling a brick by placing it at the desired location and forcing a solid connection (i.e., place-down). One robot is sufficient for these two tasks since the manipulated object is fully supported and the operation would not collapse the existing structure. On the other hand, Fig. 1 (c)-(e) showcase operations that require multi-arm collaboration. In Fig. 1 (c), the upper robot is assembling a LEGO brick on top of the character ‘S’. To establish a solid connection, the robot needs to press down and force the knob insertion due to the passive connection nature. However, since the existing connections are non-rigid, the place-down operation would break the existing overhanging structure. Therefore, the lower robot is necessary to support and stabilize the structure from below (i.e., support-bottom). Similarly, in Fig. 1 (e), the lower robot assembles a brick from the bottom onto the character ‘S’ by pushing it up and forcing the connection (i.e., place-up). The place-up operation would break the existing structure and therefore the upper robot is required to press down (i.e., support-top) in order to stabilize the structure. In addition to cooperative assembly, multi-robot collaboration also enables more dexterous object manipulation, e.g., reorienting bricks in hand. As shown in Fig. 1 (d), the upper robot initially grabs a brick from its top. To have the robot grab the brick from its bottom, we can have the upper robot handover the brick to the lower robot. With the ability to reorient objects in-hand, the robot can subsequently accomplish the assembly from the bottom as illustrated in Fig. 1 (e). Despite being a toy brand, LEGO has been widely used in entertainment, education, prototyping, etc. It is ideal for assembly benchmarking because it is a low-cost, standardized, and highly customizable assembly platform. Meanwhile, constructing LEGO structures is a challenging contact-rich manipulation problem due to the high-precision requirement and non-rigid connections. Thus, we use LEGO as our cooperative assembly benchmarking platform, and the remaining discussion is in the context of LEGO assembly.
@@
Coordinating the tasks and motions of multiple robots to accomplish cooperative assembly is challenging for several reasons. First, a system with more than a single robot introduces overhead and algorithmic challenges in their coordination. As shown in Fig. 1, certain operations involve contacts with objects (i.e., pick, place-down, place-up, handover, support-bottom, and support-top) while some do not (i.e., transit and wait). There exist delays in controlling the robot and receiving sensor feedback, or even contingencies due to unexpected events when performing contact-rich operations, which would cause delays or stops to one or more robots. The uncertainty makes it challenging to safely coordinate the robots throughout the assembly process. Second, due to the need for collaboration, robots often need to operate closely to each other as illustrated in Fig. 1 (c)-(e). Even when executing under uncertainty in a confined shared workspace, robots must always avoid collisions. Third, it is desired that the task can be accomplished more efficiently by a multi-robot system. Thus, robots must have a comprehensive understanding of the cooperative assembly task and plan optimized motions to reduce the completion time, instead of frequently stopping to avoid collisions. Lastly, the multi-robot system should scale as the complexity of assembly design increases. It is necessary for the algorithm to scale to larger and more complex tasks.
@@
Many recent works have proposed methods for planning multiple robot arms for assembly [22, 5], rearrangement [12, 14], or general manipulation tasks [45]. However, these planners often assume a sequential execution order and synchronously moving robots, where all robots start a task at the same time and wait for other robots to finish. We take a different lens and propose APEX-MR, an asynchronous planning and execution framework to coordinate multiple robots robustly, efficiently, and safely. Given an assembly task sequence, APEX-MR leverages integer-linear programming (ILP) to distribute the tasks to robots and generate a sequential robot plan. Most importantly, APEX-MR extends the temporal plan graph (TPG) [19] to multi-robot-arm systems and post-processes the sequential robot plan for asynchronous execution. Specifically, the TPG captures all dependencies between different robots’ tasks and motions and generates a partial-order graph, which can significantly reduce unnecessary wait time and is robust to execution delay and uncertainty. To highlight the applicability of our proposed algorithm within a full multi-robot assembly pipeline, we deploy the proposed APEX-MR to a dual industrial arm system to construct complex customized LEGO structures up to 258 objects in simulation and up to 47 objects in the real world. In summary, our contributions are as follows:
@@
The rest of this paper is organized as follows: Sec. II discusses relevant works. Sec. III deliberates the input that APEX-MR consumes and the assumptions we have. Sec. IV introduces our proposed APEX-MR, an asynchronous planning and execution framework for multi-robot cooperative assembly. Sec. V shows the experimental results and demonstrates the deployment to a bimanual system for LEGO assembly. Sec. VI discusses limitations and future works. Sec. VII concludes the paper.
@@ II Related Works
Multi-Agent Path Finding and Execution Multi-agent path finding (MAPF) [54] studies how to coordinate a large team of mobile robots in a discretized world environment, often in 2D grid worlds representing warehouse settings [33] or predefined roadmaps [21]. State-of-the-art MAPF algorithms can plan near-optimal collision-free paths for hundreds of robots in seconds [26].
@@
However, these MAPF algorithms achieve such impressive efficiency at the cost of neglecting robot kinematics and execution uncertainty. As a result, there is growing research within the MAPF community focused on the efficient and robust execution of (imperfect) MAPF plans on real robots. One of the most widely adopted frameworks is the temporal plan graph (TPG) [19], originally proposed to post-process MAPF plans to meet robot kinematics by enforcing passing orders at locations visited by multiple robots. TPG has since been extended to MAPF execution frameworks under uncertainty [32] and mobile robot coordination in warehouses [20, 59]. Recent advances introduce bidirectional TPG [56] and switchable TPG [1, 23], which further enhance TPG by allowing flexible passing orders at certain locations.
@@
Given the success of TPG in coordinating mobile robots, we aim to explore its applicability in coordinating robot arms. A key distinction of traditional TPG versus our use case is that the robot kinematics is more complex and may change over time as the robot arm picks up different objects, which is a key focus in this paper.
@@
Multi-Robot Arm Motion Planning Motivated by the rise of bimanual manipulation systems, many early works in the field study the problem of dual arm motion planning [51]. One naive approach that scales single-robot motion planning methods to two or more robots is to plan in the composite joint space, with methods such as RRT-Connect [25], BIT* [11], or graph of convex sets [35]. However, due to the curse of dimensionality, these methods struggle to find high-quality solutions as the number of robots increases. Other common strategies include building a composite roadmap from the Cartesian product of individual roadmaps [15], coordinating the speed of individually planned motions [2], or using prioritized planning [8]. More recently, more specialized multi-robot motion planners have been proposed that are based on roadmaps (dRRT [53], dRRT* [48], and CBS-MP [52]) or using MAPF techniques [44, 45]. Some approaches also use online planning or control techniques to generate motions in real-time. Zhang and Pecora [61] propose an online motion coordination technique, in which they plan all robots’ paths offline and a pairwise collision matrix between robots. Then, the speed of each robot can be efficiently planned online to avoid collisions, even in the presence of execution delays. However, their method cannot always find a feasible motion and relies on a task reallocation process to avoid deadlock. Other techniques such as a distributed model predictive controller [10] or a dynamical system approach [38] can also be used to generate multi-robot arm motions in real-time, but are not tailored towards long-horizon planning tasks.
@@
Multi-Robot Task and Motion Planning Beyond motion planning, multi-robot arm task and motion planning (MR-TAMP) have been studied since the 1990s, e.g., object pick and place [24], and many of which are designed for a dual arm setting [17, 12, 13, 51]. Similarly to our method, these methods take a two-stage approach in which they first generate a task plan with robot assignment and grasp poses, then search for corresponding motion plans with composite state space or prioritized planning. The authors of dRRT* have proposed extending dRRT* to multi-modal roadmaps with given possible pick-up and hand-off configurations and searching for a task and motion plan directly [46, 47]. Given that motion planning calls tend to be costly, another approach is to generate promising task plans first in a lazy manner and subsequently find corresponding motion plans and backtrack when required. This approach can be implemented with a greedy method [18], with a mixed-integer linear program [49, 5, 31] or through a satisfiability modulo theories solver [42]. However, a major drawback is that task planning often assumes motions to be synchronous, which can be suboptimal in practice. Our TPG framework is designed to be complementary to these synchronous tasks and motion planners, as it can relax the synchronicity assumption with post-processing and shorten the makespan of the overall plan. Another significant concern is that many MR-TAMP methods are designed for simple environments such as object pick and place, and their planned robot paths are executed in open loop or only in simulation. Our framework scales to more complex environments and is designed to integrate with feedback controllers and manipulation skills.
@@
Multi-Robot Arm Motion Execution Most existing work executes their multi-robot task and motion plan on real robots in a synchronized way. Since the popular motion planning framework MoveIt [7] does not natively support moving multiple robots asynchronously, some recent work has sought to address this and enable asynchronous execution. Meehan et al. [36] adds a path reservation component and treats the entire path of a moving robot arm as a static obstacle when planning another arm’s motion, which can be too conservative and cause deadlock. Stoop et al. [55] uses a central scheduler to check if a new path collides with previously scheduled paths and executes it asynchronously if there is no collision. Otherwise, the new path waits for the conflicting previous path to finish before it can start. However, their methods do not account for execution delays, may wait longer than necessary, and directly modify the MoveIt software stack. In contrast, our TPG formulation is robust to arbitrary delays by design, minimizes robot wait time, and can be implemented without modification to MoveIt.
@@
Robotic LEGO Assembly Automating LEGO assembly using robots is challenging due to the high-precision requirement, tiny sizes of LEGO bricks, and non-rigid connections in the structure. Most of the existing works address the LEGO assembly problem in simulation [43, 39], which cannot be generalized to physical assembly due to the lack of simulators to simulate the connections between LEGO bricks. Recent works [9, 16, 27, 29] assume that the structure is fully supported and only consider placing bricks on top of others, which are limited to assembling simple structures. [34] considers assembly using customized brick toys, which does not apply to LEGO assembly. In this paper, we apply APEX-MR to a bimanual system to construct complex customized LEGO structures beyond simple stacking.
@@ III Preliminaries
To coordinate a multi-robot system to perform cooperative assembly tasks, we assume that three inputs are provided.
@@
Environment Setup We assume that the environment setup, including (a) geometries of all robots, (b) poses of objects to be manipulated $\mathcal{B}=[b_{1},b_{2},\dots,b_{N_{b}}]$, and (c) states of all obstacles, is known as shown in the input section in Fig. 2. We assume that the system consists of $N$ robots and $N_{b}$ is the number of objects that can be used for the assembly. An object $b_{k}$ is semi-static, meaning it can be grasped, attached, and moved by robots. Note that since duplicate objects are common in assemblies, $b_{k}$ and $b_{k^{\prime}}$ can be identical objects, e.g., identical $2\times 2$ bricks in the character ‘S’ shown in Fig. 2.
@@
Assembly Plan Given an assembly design with $N_{a}$ objects, we assume that the assembly plan $A=[a_{1},a_{2},\dots,a_{N_{a}}]$ is provided as shown in the input section in Fig. 2. Each step $a_{j}$ refers to an object, such as a $1\times 2$ brick. The assembly plan specifies the order in which each object should be assembled. Note that due to duplicate objects in $\mathcal{B}$, there are multiple candidates $b_{i},\dots,b_{k}$ that can be used to accomplish an assembly step $a_{j}$ in $A$. Furthermore, we require the assembly plan $A$ to be physically valid [58, 30], ensuring that each task can be performed in reality. Specifically, the partially assembled structure after each step $a_{j}$ must be physically stable, and there exists at least one feasible grasp pose for each step. The assembly sequence also specifies whether each step $a_{j}$ requires two robots for cooperative assembly and the specific type of cooperation needed (i.e., support or reorientation). Lastly, we assume that $\mathcal{B}$ is sufficient so that each object required for $a_{j}$ can be found in the environment. More details on generating the assembly plan $A$ is included in Sec. V and Sec. -A.
@@
Manipulation Skills We assume that the robot manipulation skills are predefined and known. We denote the skill set as $\mathcal{S}=[s_{1},s_{2},\dots,s_{N_{s}}]$, where $N_{s}$ is the number of skills an individual robot has. For instance, a skill can be, inserting a pin, fastening a screw, picking up an object, etc. In our case of LEGO assembly, each skill is a composite of multiple motions parametrized by the pose of the manipulated object, the robot end-effector, and parameters learned from demonstrations. We assume that an algorithm, such as interpolation or RRT-Connect, can generate a reference robot path for the purpose of computing a collision-free coordination schedule. During execution, each skill is executed with a feedback controller, so that its exact execution time is unpredictable. Examples of manipulation skills are shown in Fig. 1, and more details are discussed in Sec. V and Sec. -C.
@@
In addition to these inputs, we assume the existence of a collision-free HOME pose for each robot that never blocks other moving robots from executing their tasks. Robots can also plan paths to move between different poses, or wait and hold their current pose in place, as shown in Fig. 1.